ATIC Home HASP Home LaACES Home HiSGRI Home LaSPACE Home LSU Physics
Participants Science Instrument Multimedia Status Links Private ATIC
ATIC Data Flight Results Papers ATIC Status Archive McMurdo Information

 

 


November 17, 2005

Erebus (left) and Terror (right) are both active volcanos


November 17, 2005: After close to a two and a half week delay the ATIC team was finally able to occupy Payload Bay 2 (green siding) of the new LDB facility on November 16.  Last Sunday (11/13) we were able to briefly come to the site to move ATIC into the payload bay.  This way the instrument had several days to warm up and dry out before we arrived to complete the setup.

All went very well that first day.  Besides setting up our operations area under the bay mezzanine, we also configured the electronics / mechanical area, the Silicon Matrix work area, external frame area, hodoscope / calorimeter work area and storage area.  However, the greatest triumph of the day was powering up and running ATIC in muon calibration mode.  In fact, we were able to do this within 6 hours of our initial arrival onsite.  Analysis the following day of these overnight muon runs showed that ATIC traveled very well in its custom container and only couple of hodoscope channels in S1 need to be examined more closely.

Currently, the MSFC team is in the process of updating some of the motherboards in the Silicon Matrix plus we have demounted S1 and are trying to resolve the few problems left with that detector.  If we manage to continue to avoid any major issues we may be able to move the instrument to the external framing during the first part of next week.

Several pictures from our first day of work can be found in the ATIC-3 Picture Archive.  


The new LDB facility with Mt. Discovery in the background


ATIC powered up and collecting muon calibration data on our first day of work

Track the ATIC experiment

 

   
All Flight Systems run on QNX Realtime OS.    
Site Manager:
Webmasters:
  Doug Granger
Doug Granger